OpenStudy (anonymous):

will be fan pls answer The swimming pool at Spring Valley High School is a rectangle with a base of 65 meters and a height of 25 meters. Around the perimeter of the pool is a tiled floor that extends w meters from the pool on all sides. Find an expression for the area of the tiled floor in square meters.

OpenStudy (anonymous):

Draw a rectangle for the pool and label it as given. Draw another rectangle surrounding the pool rectangle. Label the rectangle surrounding the pool as 2W +65 for the base and 2W+ 25 for the height. The area of the larger rectangle, pool plus tile, is: \[(2W+65)(2W+25)= area\]\[(2W+65)(2W+25)=4W ^{2}+180W+1625\] Now subtract the area of the pool from the area of the pool plus tile. \[4W ^{2}+180W+1625-1625=4W ^{2}+180W\] So the expression for the area of the tile is:\[4W ^{2}+180W\]